Karl's school is selling tickets to a choral performance. On the first day of ticket sales the school sold 13 senior citizen tickets and 4 child tickets for a total of $255. The school took in $255 on the second day by selling 9 senior citizen tickets and 8 child tickets. Find the price of a senior citizen and the price of a child ticket.
How would you solve this?
Thank You.

Answer by stanbon(75887)

Karl's school is selling tickets to a choral performance. On the first day of ticket sales the school sold 13 senior citizen tickets and 4 child tickets for a total of $255. The school took in $255 on the second day by selling 9 senior citizen tickets and 8 child tickets. Find the price of a senior citizen and the price of a child ticket.
--------
Value Equations:
13s + 4c = 255 dollars
9s + 8c = 255 dollars
-----
Modify for elimination::
26s + 8c = 510
9s + 8c = 255
------
19s = 255
s = $13.42 (price of a senior ticket)
-----
Solve for "c":
9s + 8c = 255
9*13.42 + 8c = 255
c = $16.78 (price of a child's ticket)
=====================
Cheers,
Stan H.
=================